OpenStudy (anonymous):

Given he following system determine conditions on a, b and c for which system will have a unique solution, an infinite number of solutions and no solutions

OpenStudy (anonymous):

x+3y-4z=8 -x-y+z=-4 2x+z=4 ax+by+cz=0

OpenStudy (anonymous):

I have examples on this question but they all have each equation set to a or b or c

OpenStudy (anonymous):

not sure how to do this I changed row 2 and three and set the first terms to zero but how do I make a 0??

OpenStudy (anonymous):

it would be a unique system if a and b and c were equal to zero, make sense? lol

OpenStudy (anonymous):

no solution if a and b and c did not equal zero

OpenStudy (anonymous):

try finding the determinant of the matrix and set it not equal to zero.the condition obtained in terms of a,b,c will then be the condition for the system to have a unique finite solution